The mean life of a tire is 30,000 km. The standard deviation is 2000 km.
a) 68% of all tires will have a life between ________ km and ________ km.
b) 95% of all tires will have a life between ________ km and ________ km.
c) What percent of the tires will have a life that exceeds 26,000 km?
d) If a company purchased 2000 tires, how many tires would you expect to last more than 28 000 km?
1
Expert's answer
2020-05-13T19:45:38-0400

Let the random variable X denotes the life of a tire and it is normally distributed: "X\sim N(\mu, \sigma^2)."

Given "\mu=30000\ km, \sigma=2000 \ km."


"X\sim N(30000,2000^2)"

(a) According to Empirical rule, the area covered between mean minus one standard deviation and mean plus one standard deviation of a normal distribution is 68%.


"P(\mu-\sigma<X<\mu+\sigma)=68\%"

"P(30000-2000<X<30000+2000)=68\%"

"P(28000<X<32000)=68\%"

68% of all tires will have a life between 28000 km and 32000 km. 


(b) According to Empirical rule, the area covered between mean minus two standard deviations and mean plus two standard deviations of a normal distribution is 95%.


"P(\mu-2\sigma<X<\mu+2\sigma)=95\%"

"P(30000-2(2000)<X<30000+2(2000))=95\%"

"P(26000<X<34000)=95\%"

95% of all tires will have a life between 26000 km and 34000 km. 


(c)


"P(X>26000)=1-P(X\leq26000)="

"=1-{1-P(26000<X<34000)\over 2}=1-{1-0.95\over 2}=0.975"

"97.5\%"


(d)


"P(X>28000)=1-P(X\leq28000)="

"=1-{1-P(28000<X<32000)\over 2}=1-{1-0.68\over 2}=0.84"


"0.84(2000)=1680"

1680 tires.